Energy (E) [M] [L] 2 /[T] 2

Distance (x) =[L]

Here [L] ` [T] and [M] denote respectively dimensions of length, time and mass


Now consider the equation. E=1/2 kx2

What must be the dimensions of k in order for the equation to be dimentally correct




1
Expert's answer
2021-10-01T13:27:23-0400

Express k in terms of other quantities:


"k=\frac{2E}{x^2},\\\space\\
k=\frac{[M] [L] ^2 [T] ^{-2}}{[L]^2}=[M][T]^{-2}."